i need help with the following please:
Write the equation of the line which has y-intercept (0, 5) and is perpendicular to the line with equation y = –3x + 1.
thank you i just do not get these

Answer by malakumar_kos@yahoo.com(315)


Write the equation of the line which has y-intercept (0, 5) and is perpendicular to the line with equation y = –3x + 1.

Let the required eq'n be y = mx+c
y - intercept = 5 i.e, c= 5
the line is perpendicular to y = -3x+1 here M = -3

the product of their slopes = -1
m*M = -1
m = -1/M = -1/-3 = 1/3
substituting for c and m we get y = (1/3)x+5